# MLLMs-Augmented Visual-Language Representation Learning

### 30 Nov 2023 v1

@@ 3.4 Text Shearing
4 Experiments
4.1 Pre-training Details
Our method is implemented in Pytorch [47] and trained on a node equipped with 8 NVIDIA A100 GPUs. For visual language pre-training, we follow the implementation of CLIP and BLIP [31] respectively. When pre-training CLIP, we use a randomly initialized ViT-B-16. For BLIP, we use a ViT-B-16 [16, 59] pre-trained on ImageNet [14] as the image encoder and $\text{Bert}_{\text{base}}$ [15] as the text encoder. Regarding image preprocessing, we employ random cropping to achieve $224\times 224$ during pre-training, and increase the image resolution to $384\times 384$ during fine-tuning. We apply RandomAugment [12] for image augmentation. For optimization, we utilize an AdamW [37] optimizer with a learning rate warm-up to 3e-4, combined with linear decay at a rate of 0.9. Our training batch size is set to 1280, and the number of training epochs is 4.
Regarding the usage of multi-modal large language models, we employed Mini-GPT4-Vicuna13B [70], Otter-Image-MPT7B  [30], Qwen-VL-Chat [5], and LLaVA-v1.5-13B [35]. During caption generation, we impose a maximum token limit of 30 to control the caption length. For the synthetic caption generation, we use beam search with number of beams equal to 1.
Table 3: The results of vision question answering, visual reasoning and image captioning of BLIP pre-trained on CC3M, CC12M and YFCC15M, respectively.
@@ 4.1 Pre-training Details
4.2 Evaluation
We evaluate the performance on the image-text retrieval task in Table 1 and the image classification task in Table 2. In addition, we provide results on more tasks in Table 3.
Image-Text retrieval. Image-text retrieval serves as a crucial metric for evaluating the bridging capability between different modalities in a model. We pre-train CLIP and BLIP on CC3M, conducting image-to-text retrieval (TR) and text-to-image retrieval (IR) testing on the MSCOCO [34] and Flickr30K [48] datasets. Results in Table 1 demonstrate significant performance improvements in both zero-shot and fine-tuning retrievals for both CLIP and BLIP. Notably, our approach enables CLIP’s zero-shot retrieval performance to even exceed that achieved by fine-tuning on the target dataset. Specifically, when using CLIP for zero-sample retrieval on MSCOCO, the R@1 of TR and IR increase by 27.2 and 19.4, respectively. For zero-shot retrieval on Flickr30K, the R@1 of TR and IR also improves by 46.1 and 35.4, respectively. Similarly, our method also achieves an improvement of 16.8$\sim$23.4 when using BLIP for zero-shot retrieval. These significant performance enhancements indicate that the pretrained model obtains improved image-text representation.
Image Classification. We assess the performance of our method on image classification task. In Table 2, we pre-train a CLIP with a VIT-B-16 architecture on CC3M and CC12M respectively, followed by testing on sixteen common image classification datasets. Our method exhibits significant performance improvements across most datasets, as demonstrated. For the CLIP pre-trained on CC3M, there’s an average improvement of 13.4 on 15 datasets and a 13.1 improvement on ImageNet [14]. For the CLIP pre-trained on CC12M, there’s an average improvement of 11.1 and 10.2. These results highlight the effective enhancement of representation learning achieved by our method.
Vision Question Answering. Visual Question Answering (VQA) [4] is the task of providing answers based on an image and a question. We conduct tests on the VQAv2 [22], A-OKVQA [53] and OK-VQA [40] using BLIP pre-trained on CC3M, CC12M and YFCC15M, respectively. The results are presented in Table 3. Similar to BLIP [31], we also consider VQA as an answer generation task. The consistent performance improvements obtained through our method indicate that the model has acquired more robust visual-language representation from datasets enriched with MLLMs’ knowledge. Besides, the improvement on A-OKVQA and OK-VQA indicate that the model has more common sense and world knowledge. These results suggest an enhanced capability in visual understanding and language capabilities together.
@@ 4.2 Evaluation
Table 4: The results of performance on the video-language retrieval dataset MSRVTT.
Image Caption Generation. Image Captioning is the task of generating a text description of the image content given an image. We conduct tests on COCO and Nocaps [2] using the BLIP pre-trained on CC3M, CC12M and YFCC15M, respectively. Following the approach in [31] , we initially fine-tune the pre-trained model with the LM loss on COCO. To achieve better results, we also add ”a picture of” at the beginning of the prompt. The results are presented in Table 3, indicating improvements in BLEU@4 and CIDEr metrics. These findings suggest that our method enhances the model’s understanding of the relationship between text and images, resulting in higher similarity with human annotations during the captioning process.
@@
Figure 6: The influence of the token length, batch size, model numbers and epochs on visual-language pre-training
Video-Language Task. Text-to-video retrieval is a metric for evaluating a model’s generalization ability in video-language tasks. We evaluate the effectiveness of our method on the MSRVTT dataset. Following the  [31], we fine-tune the model on COCO. For video input, we uniformly sample 8 frames from it to get a sequence. The results in the Table 3 demonstrate a stable performance improvement achieved by our method. This suggests that good visual-language representation learning may be key to video-text retrieval.
4.3 Analysis
Notably, as shown in Table 3, our method shows certain improvements when pre-training BLIP on CC12M [8] and YFCC15M [58]. This indicates that our method is scalable to larger datasets. It also emphasizes the significance of constructing well-represented image-text pairs for enhancing visual-language pre-training.
@@ 4.3 Analysis
Table 5: Retrieval performance of models trained on captions rewritten using one MLLM.
4.4 Ablation Study
Figure 7: Long captions tend to have higher similarity due to the style limitations of the model itself. Similar content in captions is shown in red.
Caption Length. The length of the generated captions is one of the factors affecting visual-language pre-training. We utilize MiniGPT-4 to generate captions for CC3M with varying max token number limits. By training BLIP on captions of different lengths and conducting retrieval on MSCOCO , we present the results in Fig.5(a) and observe that as the length of the caption increases, the model’s performance tends to decrease.
This may be due to the fact that too long captions cause more similar text features between captions of different images. We illustrate this phenomenon in Fig. 7. When the caption length increases, the generated captions often become identical in many instances. This leads to different images easily mapping to the same text features, creating challenges in learning an effective representation.
Batch size. We explore the impact of batch size on training in the Fig. 5(b). Training models on synthetic captions requires relatively high batch sizes compared to using real captions annotated by humans. We attribute this to the fact that, with small batch size, the main gradient information obtained by the model is noisy, differing significantly from the true gradient. Conversely, with an increased batch size, the gradient of the sampled distribution becomes more similar to the gradient of the real distribution, enabling the model to generalize better.
Number of MLLMs. We explore the impact of the number of MLLMs on the transfer performance in the Fig. 5(c). We add four MLLMs{MiniGPT-4, Otter, Qwen-VL, and LLAVA-1.5}in order. Using pre-trained models for zero-shot retrieval on MSCOCO, we observe that, as the number of MLLMs increases, the performance gradually improves. However, the magnitude of this improvement is gradually decreasing, indicating that the one image’s caption information obtained from MLLM is reaching saturation.
Training epochs. We visualize the curve depicting the number of training epochs versus the model’s performance on the retrieval task in the Fig. 5(d). It is evident that our method achieves promising performance with only a small number of epochs. In the main experiment, we choose the epoch number of 4 to compare with the baseline without increasing the training cost. When further increasing the number of epochs, our method continues to show certain improvements.
Comparison using only one MLLM. We test the rewriting of captions using only one MLLM and utilize the generated captions for training. Captions are rewritten using MiniGPT-4, Otter, Qwen-VL, and LLaVA-1.5, and the results are presented in the Table 5. The results indicates that there is an upper limit to the performance improvement achievable by rewriting with only one model. Combining multiple MLLMs proves to be more effective in enhancing visual-language representation learning.
Figure 9: Cosine similarity distribution of images and text. Our method improves the average matching degree of image-text pairs.
Figure 10: Distribution of images and text. Our method aligns images and text better than the original distribution.
@@ 4.5 Visualization
Image and text distribution visualization. We visualize the feature distribution of image-text pairs from the original CC3M and ours dataset in the Fig. 10. It can be observed that the original CC3M dataset contains numerous discrete and unmatched image-text pairs. In contrast, our dataset exhibits a distribution where almost all images have corresponding texts. This distribution perspective explains how our method aligns more continuous images with discrete language, thereby enhancing visual-language representation learning.
Image captioning visualization. In the Fig. 8 we illustrate the difference in image captioning between models pre-trained using the original CC3M dataset and our improved dataset. without any fine-tuning, our model exhibits a significant improvement in the ability to recognize concepts in images. We provide more examples of image captioning in the appendix.
Cosine similarity visualization. We visualize the cosine similarity comparison between the original CC3M dataset and our dataset in the Fig. 9. We use a pre-trained CLIP to calculate the cosine distance between images and texts. It is evident that our dataset exhibits a higher average similarity than the original distribution.
@@ 5 Conclusion and Discussion
In this paper, we propose to augment visual-language representation learning by leveraging MLLMs. While retaining the rich visual information of the original dataset, we utilize multiple MLLMs to extend captions for each image. Additionally, we introduce “text shearing” to address issues of hallucinations and intrinsic caption style in synthetic captions. Validated across various VLP frameworks and datasets, our method significantly improves performance on numerous downstream tasks. This encourages further exploration in the utilization of MLLMs in future.
Limitations and future works Although our method has achieved remarkable results in enhancing visual-language representation learning, a certain proportion of noise still persists due to unreliable MLLM’s outputs. These noises limit further improvement of model’s performance. Future research could explore leveraging more powerful MLLMs and expanding to larger datasets.

### 13 Mar 2024 v3

@@ 4.2 Evaluation
Image Classification. We evaluate the performance of our method on the task of image classification. In Table 2, we pre-train a CLIP with a VIT-B-16 architecture on CC3M and CC12M respectively, followed by testing on sixteen common image classification datasets. We compare our results with  [17]. Our method exhibits significant performance improvements across most datasets, as demonstrated. For the CLIP pre-trained on CC3M, there’s an average improvement of 13.4 on 15 datasets and a 13.1 improvement on ImageNet [14]. For the CLIP pre-trained on CC12M, there’s an average improvement of 11.1 and 10.2 respectively. These results highlight the effective enhancement of representation learning achieved by our method.
Table 3: The linear probing results of our method on 15 common classification datasets with CLIP pre-trained on CC3M and CC12M. On CC3M, our method demonstrates an average improvement of 7.9 on 15 datasets.
Data Model
Food-101
CIFAR-10
CIFAR-100
SUN397
Cars
Aircraft
DTD
Pets
Caltech-101
Flowers
STL-10
EuroSAT
RESISC45
GTSRB
Country211
Average
$\Delta$(%)
Model Architecture: ViT-B/16
CC3M CLIP 62.6 86.8 68.1 58.5 32.8 40.9 63.4 69.6 82.0 89.4 91.7 95.9 89.0 71.9 13.3 67.7 -
LaCLIP [17] 63.8 87.7 69.5 60.2 32.4 42.7 64.0 71.1 83.3 90.2 93.4 95.8 89.7 74.6 13.2 68.8 $\uparrow$1.1
CLIP 52.8 79.7 58.8 52.4 21.9 26.4 53.1 56.6 77.5 76.6 85.7 93.9 84.5 66.7 8.7 59.7 -
Ours 64.0 87.7 68.5 59.1 34.5 32.1 60.4 73.3 85.5 83.6 92.6 95.3 87.9 78.4 10.6 67.6 $\uparrow$ 7.9
CC12M CLIP 81.6 93.8 79.3 72.0 75.1 52.6 75.6 86.2 92.2 95.3 97.3 96.7 93.1 80.6 19.7 79.4 -
LaCLIP [17] 82.9 94.7 79.7 73.8 79.9 54.5 75.7 87.7 93.0 96.4 98.0 96.4 93.0 81.9 19.7 80.5 $\uparrow$1.1
CLIP 73.9 89.9 71.2 68.9 62.8 35.6 67.1 81.1 90.3 87.7 95.0 95.6 89.1 77.4 13.8 73.3 -
Ours 80.6 93.9 78.1 72.1 62.1 35.6 72.1 82.7 90.6 87.9 97.3 95.7 91.1 84.8 13.8 75.9 $\uparrow$ 2.6
Linear-Probing. We also explore the linear-probing performance in image classification. We use CLIP pre-trained on CC3M and CC12M for evaluation respectively. Similar to the zero-shot image classification, we compare our method with [17] across 15 common datasets. The results in Table 3 show that our method also has certain improvements for linear probing. This shows that rich text concepts contribute to the effective training of visual encoders. As the visual encoder learns each image, it benefits from supervision through multiple accurate and diverse captions. This allows it to efficiently acquire a universal representation of the image, leading to a more effective characterization of the image and text embedding space.
Vision Question Answering. Visual Question Answering (VQA) [4] is the task of providing answers based on an image and a question. We conduct tests on the VQAv2 [22], A-OKVQA [53] and OK-VQA [40] using BLIP pre-trained on CC3M, CC12M and YFCC15M, respectively. The results are presented in Table 4. Similar to BLIP [32], we also consider VQA as an answer-generation task. The consistent performance improvements obtained through our method indicate that the model has acquired a more robust visual-language representation from datasets enriched with MLLMs’ knowledge. Besides, the improvement on A-OKVQA and OK-VQA indicates that the model has more common sense and world knowledge. These results suggest an enhanced capability in visual understanding and language capabilities together.
Table 4: The results of vision question answering, visual reasoning and image captioning of BLIP pre-trained on CC3M, CC12M, and YFCC15M, respectively.
@@
Image Captioning. Image Captioning is the task of generating a text description of the image content given an image. We conduct tests on COCO and Nocaps [2] using the BLIP pre-trained on CC3M, CC12M, and YFCC15M, respectively. Following the approach in [32], we initially fine-tune the pre-trained model with the LM loss on COCO. To achieve better results, we also add "a picture of" at the beginning of the prompt. The results are presented in Table 4, indicating improvements in BLEU@4 and CIDEr metrics. These findings suggest that our method enhances the model’s understanding of the relationship between texts and images, resulting in higher similarity with human annotations during the captioning process.
@@
Table 5: The retrieval performance on the video-language retrieval dataset MSRVTT. R@K stands for recall at K and MdR stands for median rank.
Video-Language Task. Text-to-video retrieval is a metric for evaluating a model’s generalization ability in video-language tasks. We evaluate our method on the MSRVTT dataset. Following the  [32], we fine-tune the model on COCO. For video input, we uniformly sample 8 frames from it to get a sequence. The results in Table 4 demonstrate a stable performance improvement achieved by our method. This suggests that robust visual-language representation learning may be key to video-text retrieval.
4.3 Analysis
Scaling Ability. Notably, as shown in Table 4, our method shows certain improvements when pre-training BLIP on CC12M [8] and YFCC15M [58]. This indicates that our method is scalable to larger datasets to some extent. It also emphasizes the significance of constructing well-represented image-text pairs for enhancing visual-language pre-training.
Training Cost. To ensure a fair comparison with baseline methods, we carefully craft the training schedule. Considering that the number of image-text pairs in the augmented dataset is $k$ times the number of image-text pairs in the original dataset, we adjust the pre-training epochs to be $1/k$ of the original epochs. By doing this, we avoid introducing additional training overhead.
@@ 4.3 Analysis
Figure 5: The influence of the token length, batch size, model numbers, and epochs on the performance of visual-language pre-training.
4.4 Ablation Study
Caption Length. The length of the generated captions is one of the factors affecting visual-language pre-training. We utilize MiniGPT-4 to generate captions for CC3M with varying max token number limits. By training BLIP on captions of different lengths and conducting retrieval on MSCOCO, we present the results in Figure4(a) and observe that as the length of the caption increases, the model’s performance tends to decrease. This may be because too-long captions
Figure 6: Long captions for different images tend to be similar due to the intrinsic inference manner of the model. Similar content in captions is shown in red.
cause more similar text features between captions of different images. We illustrate this phenomenon in Figure 6. When the caption length increases, the generated captions often become identical in many instances. This leads to different images easily mapping to the same text features, creating challenges in learning an effective representation.
Batch Size. We explore the impact of batch size on training in Figure 4(b). Training models on synthetic captions requires a relatively high batch size compared to using captions annotated by humans. We attribute this to the fact that, with a small batch size, the main gradient information obtained is noisy, differing significantly from the true gradient. Conversely, with an increased batch size, the gradient of the sampled distribution becomes more similar to the gradient of the real distribution, enabling the model to generalize better.
Number of MLLMs. We explore the impact of the number of MLLMs on the transfer performance in Figure 4(c). We add four MLLMs{MiniGPT-4, Otter, Qwen-VL, and LLAVA-1.5}in order. Using pre-trained models for zero-shot retrieval on MSCOCO, we observe that, as the number of MLLMs increases, the performance gradually improves. However, the magnitude of this improvement is gradually decreasing, indicating that the one image’s caption information obtained from MLLM is reaching saturation.
Training Epochs. We visualize the curve depicting the number of training epochs versus the model’s performance on the retrieval task in Figure 4(d). It is evident that our method achieves promising performance with only a small number of epochs. In the main experiment, we choose epoch number 4 to compare with the baseline without increasing the training cost. When further increasing the number of epochs, our method continues to show certain improvements.
@@ 4.4 Ablation Study
Table 6: The retrieval performance of models trained on captions rewritten using one MLLM.
MLLM’s Independent Effect. We also evaluate the performance of rewriting captions using only one MLLM and utilize the generated captions for training. Captions are rewritten using MiniGPT-4, Otter, Qwen-VL, and LLaVA-1.5 respectively, and the results are presented in Table 6. The results indicate that there is an upper limit to the performance improvement achievable by rewriting with only one MLLM. Besides, different MLLMs also have different strengths in knowledge. Combining multiple MLLMs’ knowledge proves to be more effective in enhancing visual-language pre-training.
@@ 4.5 Visualization
Image Captioning Visualization. In Figure 7 we illustrate the difference in image captioning between models pre-trained on the original CC3M dataset and our improved dataset. Without any fine-tuning, our model exhibits a significant improvement in the ability to recognize visual concepts in images.
(a) Image-text distribution visualization.
(b) Cosine similarity visualization.
Figure 8: (a) Distribution of images and texts in the original CC3M dataset and our enhanced CC3M dataset. Our method aligns images and texts better in the latent space. (b) Cosine similarity distribution of images and texts. Our method improves the average matching degree of image-text pairs.
Image-text Distribution Visualization. We visualize the feature distribution of image-text pairs from the original CC3M and our dataset in Figure 7(a). It can be observed that the original CC3M dataset contains numerous discrete and unmatched image-text pairs. In contrast, our dataset exhibits a distribution where almost all images have corresponding texts. This distribution perspective explains how our method aligns more continuous images with discrete language, thereby enhancing visual-language representation learning.
Cosine Similarity Visualization. We visualize the cosine similarity comparison between the original CC3M dataset and our enhanced CC3M dataset in Figure 7(b). We use a pre-trained CLIP to calculate the cosine distance between image embeddings and text embeddings. Our method exhibits a higher average similarity, which proves the effectiveness in improving the matching degree of the image-text pairs and establishing robust image-text associations for datasets.
4.6 Comparison with More Methods
Table 7: The image-text retrieval results on COCO and Flickr30K and classification results on ImageNet and ImageNetV2 [50] using the pre-trained CLIP on CC3M.
Method  COCO (R@1) Flickr30k (R@1) ImageNet ImageNetV2 Avg.
I2T T2I I2T T2I
CLIP 13.9 9.6 26.3 18.0 14.6 12.5 15.8
VeCLIP [29] 32.0 22.1 57.2 36.5 20.7 17.9 31.1
CLIP 8.7 7.1 17.4 13.9 11.9 10.3 11.6
Ours 35.9 26.5 63.5 49.3 25.0 21.4 36.9
We also compare our work with the contemporaneous work VeCLIP [29] in the same setting in Table 7. We use pre-trained CLIP to compare results on image-text retrieval and image classification. It can be seen that our method achieves better performance improvement than the caption fusion used in VeCLIP. This also implies that establishing more continuous image-text correspondences is the key to improving visual-language pre-training on small-scale datasets. We also compare our method with  [42] in the appendix.
@@ 5 Conclusion and Discussion
In this paper, we propose to augment visual-language representation learning by leveraging multiple MLLMs. While retaining the rich visual information of the original dataset, we utilize multiple MLLMs to extend diverse captions for each image. Additionally, we introduce “text shearing" to address issues of hallucinations and monotonous language style in synthetic captions. Validated across various visual-language pre-training frameworks and datasets, our method significantly improves performance on numerous downstream tasks. This encourages further exploration in the utilization of MLLMs in the future.
Limitations and future works Although our method has achieved remarkable results in enhancing visual-language representation learning, a certain proportion of noise persists due to unreliable MLLMs’ outputs. These noises limit the further improvement of the model’s performance. Future research could explore leveraging more powerful MLLMs to generate accurate captions and expanding to larger datasets.